assume that the population consists of 7 containers having the following weight 6,7,8,9,10,11,12. Find mean μ and standard deviation.


1
Expert's answer
2021-12-24T08:57:27-0500
"mean=\mu=\dfrac{6+7+8+9+10+11+12}{7}=9"

"\sigma^2=\dfrac{1}{7}((6-9)^2+(7-9)^2+(8-9)^2"

"+(9-9)^2+(10-9)^2+(11-9)^2+(12-9)^2)"

"=4"

"\sigma=\sqrt{\sigma^2}=\sqrt{4}=2"

Mean "\mu=9," standard deviation "\sigma=2."